# EWMBench: Evaluating Scene, Motion, and Semantic Quality in Embodied World Models

### v1

@@ 1 Introduction
Despite advancements in EWMs, a fundamental question remains unresolved: “How can we determine whether a video generation model qualifies as a good embodied world model, beyond merely serving as a general-purpose video generator?” Addressing this question is essential for guiding model development and assessing a model’s ability to produce physically grounded, action-consistent behaviors. While existing video generation benchmarks Huang et al. (2024a) focus on perceptual metrics like visual fidelity, language alignment, and human preference, these criteria are insufficient for evaluating EWMs. Embodied generation tasks have unique requirements, such as coherence in embodiment motion and plausibility in action execution, as illustrated in Figure 1. For instance, in robotic manipulation scenarios, the background, object configuration, and embodiment structure (e.g., robot morphology) are expected to remain static, while only the robot’s pose and interactions evolve according to instructions. This structured realism sets EWMs apart from general video generators and demands evaluation beyond conventional criteria.
@@
To benchmark EWMs under our proposed criteria, we construct a comprehensive colosseum consisting of a curated benchmark dataset and open-source evaluation tools. The dataset is built on Agibot-World AgiBot (2024), the largest real-world robotic manipulation dataset, featuring diverse tasks at scale. We select 30 candidate samples across ten tasks with clear action-ordering constraints, where correct execution requires understanding logical dependencies and affordances—posing significant challenges for embodied video generation. For each sample, static initial frames are clipped to ensure subsequent frames strictly reflect annotated language instructions without redundant movements. To reflect task diversity, we account for cases where multiple trajectories achieve the same goal and incorporate voxelized scoring to encourage variation. With this dataset, initial frames and language instructions are fed into different video generators, and the generated videos are compared against ground truth (GT) using various metrics.
@@ 2.1 Video Generative Models
Diffusion-based video generation models have made significant advances in recent years, particularly in text-to-image (T2I) generation Blattmann et al. (2023); Ho et al. (2020); Song et al. (2020)and text-to-video (T2V)  Zhang et al. (2023); Chen et al. (2023); Ren et al. (2024); Guo et al. (2023); Xing et al. (2025) generation. Recent works Kong et al. (2024); Zheng et al. (2024); Bao et al. (2024) have explored the usage of Diffusion Transformers in the denoising process. Some researchers Cheang et al. (2024) in robot learning aim to leverage this knowledge to address the challenge of data scarcity in robotic data collection. In some works Brooks et al. (2024); Yang et al. (2023)video models are employed to predict and simulate future states of dynamic systems. In robotics, video models have been utilized to predict future frames based on textual and visual inputs Zhou et al. (2024); Huang et al. (2025); Chi et al. (2024).
@@ 2.2 Evaluation of Video Generative Models
TC-Bench Feng et al. (2024)
Physics-IQ Motamed et al. (2025)
VBench Huang et al. (2024a)
VBench++ Huang et al. (2024b)
PhyGenBench Meng et al. (2024)
T2V-CompBench Sun et al. (2024)
VMBench Ling et al. (2025)
EvalCrafter Liu et al. (2024)
T2VBench Ji et al. (2024)
EVA Chi et al. (2024)
@@
With the rapid development of video generation models and their broad applications across various domains, evaluating these models has become increasingly important. Earlier approaches primarily relied on conventional metrics such as Fréchet Inception Distance (FID)Heusel et al. (2017), Inception Score (IS)Salimans et al. (2016), and Fréchet Video Distance (FVD) Unterthiner et al. (2019). However, these metrics primarily focus on visual appearance and offer limited insights into the diverse and complex capabilities of modern video generation models. Recent evaluation frameworks Huang et al. (2024a, b) introduced a more structured approach that considers multiple capability dimensions. Nevertheless, these benchmarks still emphasize visual quality, including aesthetic appeal and motion smoothness. To address these limitations, specialized benchmarks have emerged. For instance, PhyGenBench Meng et al. (2024) evaluates a model’s understanding of physical laws using Vision-Language Models (VLMs), while T2V-CompBench Sun et al. (2024) assesses compositionality, covering motion, actions, spatial relationships, and attributes.
@@ 4 Experiments
Models We evaluate EWMBench across seven video generation models categorized as open-source, commercial, and domain-adapted. The open-source models include OpenSora 2.0Peng et al. (2025), which demonstrates strong performance on VBench with low training costs; LTXHaCohen et al. (2024), capable of real-time generation for interactive tasks; and COSMOS-7BAgarwal et al. (2025), pre-trained for digital twin applications. Commercial models optimized for zero-shot generation include Kling-1.6Kuaishou (2025) and Hailuo I2V-01-liveHailuo (2025), both of which rank highly in recent benchmarks. Domain-adapted models fine-tuned for embodied scene understanding and action prediction include LTX_FT, a fine-tuned version of LTX, and EnerVerse, which is specifically designed for embodied scenarios. At present, we primarily focus on the Image-Text-to-Video setting, as no open-source action-conditioned video generation models are currently available. The evaluation of such models is left for future work. Nonetheless, with our unified input format and visual-space evaluation operations, EWMBench could also support this evaluation in the future. For the current evaluation, we tested 10 tasks, each consisting of 10 ground-truth episodes. Three videos were generated per model per episode, and the best prediction was selected using a best-of-three strategy, resulting in a total of 2,100 videos.
@@ 4.1 Evaluation Results
Comm. KlingKuaishou (2025) 0.8888 0.3231 0.3047 0.3162 0.9440 0.0493 0.1675 0.8535 0.9667 2.0370 3.8698
HailuoHailuo (2025) 0.8577 0.2229 0.1344 0.1789 0.5362 0.0370 0.1848 0.8857 0.9111 2.0186 3.4125
Open. COSMOSAgarwal et al. (2025) 0.7963 0.2500 0.2052 0.2533 0.7085 0.0803 0.1230 0.8458 0.7333 1.7824 3.2872
OpenSoraPeng et al. (2025) 0.9210 0.1548 0.0474 0.1420 0.3442 0.0415 0.1598 0.8505 0.8222 1.8739 3.1392
LTXHaCohen et al. (2024) 0.9156 0.1575 0.1002 0.1425 0.4002 0.0174 0.0687 0.8324 0.7333 1.6518 2.9676
@@ 4.2 Human Evaluation
We then compared this aggregated human rankings (Figure 6 (A)) with those produced by EWMBench and VBench Huang et al. (2024a), one of the most popular video generation evaluation benchmarks. As shown in Figure 6 (B), EWMBench’s rankings align more closely with human judgments than VBench rankings, indicating stronger consistency with human perception.
@@ A.1 Additional Details on World Specification
Trajectory Extraction We extract 2D trajectories of both end-effectors using a consistent detection pipeline. Specifically, we apply a fine-tuned YOLO-World Cheng et al. (2024) model for per-frame detection and use BoT-SORT  Aharon et al. (2022)for temporal association. To ensure fair evaluation, we compute the convex hull of each hand’s trajectory and select the hand with the largest spatial extent—measured by the maximum Euclidean distance between any two points on the hull—as the primary trajectory for motion evaluation.
@@ A.2.1 Task Selection and Scene Diversity
To support evaluation across a wide range of manipulation scenarios, we select 10 representative tasks from the Agibot-World AgiBot (2024) dataset. The selection prioritizes tasks with clear operational goals, sequential dependencies, and diverse object interactions. These tasks span both household and industrial contexts and are designed to challenge models across spatial reasoning, tool-use, and action ordering.
@@ A.2.2 Action Motion Sampling Strategy
To promote trajectory diversity in the final evaluation subset, we adopt a greedy selection algorithm Vince (2002):
@@ A.3 Additional Details on Metrics
Fine-tuning DINOv2 for Embodied Scenes. To enhance the extraction of task-relevant visual features in embodied operation scenarios, we fine-tune DINOv2 Oquab et al. (2023) on the Agibot-World dataset using 20,000 iterations of unsupervised training. We use the dinov2-vitb14-reg4 checkpoint as the initialization. We visualize feature maps from three model variants: (1) the original DINO Zhang et al. (2022) (ViT-B/16, used in VBench), (2) pre-trained DINOv2, and (3) our fine-tuned DINOv2.
@@ A.3.2 Action Motion Metrics
The Symmetric Hausdorff DistanceSerra (1998) (SymH) measures the maximum spatial deviation between the generated trajectory and the corresponding ground truth trajectory. This distance represents the greatest of the minimum distances between points from both trajectories. HSD is particularly useful for evaluating the spatial alignment of the generated trajectory with the true trajectory, ensuring that the generated path adheres to expected movement patterns and does not deviate significantly from the true action trajectory.
@@
NDTWMüller (2007) is used to evaluate the overall shape similarity and temporal alignment of trajectories. While HSD focuses on spatial deviations, NDTW evaluates the overall trajectory shape and how well the generated actions align with the timing and sequence of the true actions. This metric is particularly useful for capturing the temporal causality and task sequencing that the model should learn from the true trajectory.
@@
Velocity and acceleration are critical components in robotic control, directly affecting the physical feasibility of the generated actions. To evaluate how well the predicted trajectories align with the real-world motion characteristics, we extract the 2D velocity and acceleration time series from both predicted and ground-truth trajectories. We then compute the Wasserstein distanceVillani and Villani (2009) (Earth Mover’s Distance, EMD) between these distributions to quantify their differences.

### v2

@@ 1 Introduction
Despite advancements in EWMs, a fundamental question remains unresolved: “How can we determine whether a video generation model qualifies as a good embodied world model, beyond merely serving as a general-purpose video generator?” Addressing this question is essential for guiding model development and assessing a model’s ability to produce physically grounded, action-consistent behaviors. While existing video generation benchmarks [18] focus on perceptual metrics like visual fidelity, language alignment, and human preference, these criteria are insufficient for evaluating EWMs. Embodied generation tasks have unique requirements, such as coherence in embodiment motion and plausibility in action execution, as illustrated in Figure 1. For instance, in robotic manipulation scenarios, the background, object configuration, and embodiment structure (e.g., robot morphology) are expected to remain static, while only the robot’s pose and interactions evolve according to instructions. This structured realism sets EWMs apart from general video generators and demands evaluation beyond conventional criteria.
@@
To benchmark EWMs under our proposed criteria, we construct a comprehensive colosseum consisting of a curated benchmark dataset and open-source evaluation tools. The dataset is built on Agibot-World [2], the largest real-world robotic manipulation dataset, featuring diverse tasks at scale. We select 30 candidate samples across ten tasks with clear action-ordering constraints, where correct execution requires understanding logical dependencies and affordances—posing significant challenges for embodied video generation. For each sample, static initial frames are clipped to ensure subsequent frames strictly reflect annotated language instructions without redundant movements. To reflect task diversity, we account for cases where multiple trajectories achieve the same goal and incorporate voxelized scoring to encourage variation. With this dataset, initial frames and language instructions are fed into different video generators, and the generated videos are compared against ground truth (GT) using various metrics.
@@ 2.1 Video Generative Models
Diffusion-based video generation models have made significant advances in recent years, particularly in text-to-image (T2I) generation [5, 16, 33]and text-to-video (T2V)  [41, 8, 30, 12, 38] generation. Recent works [21, 42, 4] have explored the usage of Diffusion Transformers in the denoising process. Some researchers [7] in robot learning aim to leverage this knowledge to address the challenge of data scarcity in robotic data collection. In some works [6, 39]video models are employed to predict and simulate future states of dynamic systems. In robotics, video models have been utilized to predict future frames based on textual and visual inputs [43, 17, 10].
@@ 2.2 Evaluation of Video Generative Models
TC-Bench [11]
Physics-IQ [26]
VBench [18]
VBench++ [19]
PhyGenBench [25]
T2V-CompBench [34]
VMBench [23]
EvalCrafter [24]
T2VBench [20]
EVA [10]
@@
With the rapid development of video generation models and their broad applications across various domains, evaluating these models has become increasingly important. Earlier approaches primarily relied on conventional metrics such as Fréchet Inception Distance (FID)[15], Inception Score (IS)[31], and Fréchet Video Distance (FVD) [35]. However, these metrics primarily focus on visual appearance and offer limited insights into the diverse and complex capabilities of modern video generation models. Recent evaluation frameworks [18, 19] introduced a more structured approach that considers multiple capability dimensions. Nevertheless, these benchmarks still emphasize visual quality, including aesthetic appeal and motion smoothness. To address these limitations, specialized benchmarks have emerged. For instance, PhyGenBench [25] evaluates a model’s understanding of physical laws using Vision-Language Models (VLMs), while T2V-CompBench [34] assesses compositionality, covering motion, actions, spatial relationships, and attributes.
@@ 4 Experiments
Models We evaluate EWMBench across seven video generation models categorized as open-source, commercial, and domain-adapted. The open-source models include OpenSora 2.0[29], which demonstrates strong performance on VBench with low training costs; LTX[13], capable of real-time generation for interactive tasks; and COSMOS-7B[1], pre-trained for digital twin applications. Commercial models optimized for zero-shot generation include Kling-1.6[22] and Hailuo I2V-01-live[14], both of which rank highly in recent benchmarks. Domain-adapted models fine-tuned for embodied scene understanding and action prediction include LTX_FT, a fine-tuned version of LTX, and EnerVerse, which is specifically designed for embodied scenarios. At present, we primarily focus on the Image-Text-to-Video setting, as no open-source action-conditioned video generation models are currently available. The evaluation of such models is left for future work. Nonetheless, with our unified input format and visual-space evaluation operations, EWMBench could also support this evaluation in the future. For the current evaluation, we tested 10 tasks, each consisting of 10 ground-truth episodes. Three videos were generated per model per episode, and the best prediction was selected using a best-of-three strategy, resulting in a total of 2,100 videos.
@@ 4.1 Evaluation Results
Comm. Kling[22] 0.8888 0.3231 0.3047 0.3162 0.9440 0.0493 0.1675 0.8535 0.9667 2.0370 3.8698
Hailuo[14] 0.8577 0.2229 0.1344 0.1789 0.5362 0.0370 0.1848 0.8857 0.9111 2.0186 3.4125
Open. COSMOS[1] 0.7963 0.2500 0.2052 0.2533 0.7085 0.0803 0.1230 0.8458 0.7333 1.7824 3.2872
OpenSora[29] 0.9210 0.1548 0.0474 0.1420 0.3442 0.0415 0.1598 0.8505 0.8222 1.8739 3.1392
LTX[13] 0.9156 0.1575 0.1002 0.1425 0.4002 0.0174 0.0687 0.8324 0.7333 1.6518 2.9676
@@ 4.2 Human Evaluation
We then compared this aggregated human rankings (Figure 6 (A)) with those produced by EWMBench and VBench [18], one of the most popular video generation evaluation benchmarks. As shown in Figure 6 (B), EWMBench’s rankings align more closely with human judgments than VBench rankings, indicating stronger consistency with human perception.
@@ A.1 Additional Details on World Specification
Trajectory Extraction We extract 2D trajectories of both end-effectors using a consistent detection pipeline. Specifically, we apply a fine-tuned YOLO-World [9] model for per-frame detection and use BoT-SORT  [3]for temporal association. To ensure fair evaluation, we compute the convex hull of each hand’s trajectory and select the hand with the largest spatial extent—measured by the maximum Euclidean distance between any two points on the hull—as the primary trajectory for motion evaluation.
@@ A.2.1 Task Selection and Scene Diversity
To support evaluation across a wide range of manipulation scenarios, we select 10 representative tasks from the Agibot-World [2] dataset. The selection prioritizes tasks with clear operational goals, sequential dependencies, and diverse object interactions. These tasks span both household and industrial contexts and are designed to challenge models across spatial reasoning, tool-use, and action ordering.
@@ A.2.2 Action Motion Sampling Strategy
To promote trajectory diversity in the final evaluation subset, we adopt a greedy selection algorithm [37]:
@@ A.3 Additional Details on Metrics
Fine-tuning DINOv2 for Embodied Scenes. To enhance the extraction of task-relevant visual features in embodied operation scenarios, we fine-tune DINOv2 [28] on the Agibot-World dataset using 20,000 iterations of unsupervised training. We use the dinov2-vitb14-reg4 checkpoint as the initialization. We visualize feature maps from three model variants: (1) the original DINO [40] (ViT-B/16, used in VBench), (2) pre-trained DINOv2, and (3) our fine-tuned DINOv2.
@@ A.3.2 Action Motion Metrics
The Symmetric Hausdorff Distance[32] (SymH) measures the maximum spatial deviation between the generated trajectory and the corresponding ground truth trajectory. This distance represents the greatest of the minimum distances between points from both trajectories. HSD is particularly useful for evaluating the spatial alignment of the generated trajectory with the true trajectory, ensuring that the generated path adheres to expected movement patterns and does not deviate significantly from the true action trajectory.
@@
NDTW[27] is used to evaluate the overall shape similarity and temporal alignment of trajectories. While HSD focuses on spatial deviations, NDTW evaluates the overall trajectory shape and how well the generated actions align with the timing and sequence of the true actions. This metric is particularly useful for capturing the temporal causality and task sequencing that the model should learn from the true trajectory.
@@
Velocity and acceleration are critical components in robotic control, directly affecting the physical feasibility of the generated actions. To evaluate how well the predicted trajectories align with the real-world motion characteristics, we extract the 2D velocity and acceleration time series from both predicted and ground-truth trajectories. We then compute the Wasserstein distance[36] (Earth Mover’s Distance, EMD) between these distributions to quantify their differences.